OpenStudy (anonymous):

Hi - I've one equation left but just cannot solution it. (9^y)/(27^(2y+1))=81 and I know that y=-7/4 but I just can't get the working right - any help/guidance is appreciated - thanks. Kev.

OpenStudy (anonymous):

9=3^2 27=3^3

OpenStudy (anonymous):

write everything with the same base use 3

OpenStudy (anonymous):

what fiddle said. get \[\frac{3^{2y}}{3^{3(2y+1)}}=3^4\]

OpenStudy (anonymous):

I was working on that - and got it to 3^2y/3^6y+1 = 3^4 and then knocking the base out to give me 2y/6y+1 = 4 but that seemed to easy

OpenStudy (anonymous):

ahh, yep as satellite73 has listed (thanks, I'm on the right track) but now I just don't seem able to balance the thing :(

OpenStudy (anonymous):

because it is wrong. you have so subtract the exponents

OpenStudy (anonymous):

\[2y-3(2y+1)=4\]

OpenStudy (anonymous):

that works out to -7/4

OpenStudy (anonymous):

and don't forget the distributive property

OpenStudy (anonymous):

\[2y-6y-3=4\] etc

OpenStudy (anonymous):

face+palm = doh! Only been on this one for 2hrs and just couldn't get it

OpenStudy (anonymous):

thanks folks - feel like such a dork. It's going to teach me to offer to help the kids with their A level Maths homework - I managed the other five they couldn't get - but this one was beyond my teaching of 20yrs ago.

OpenStudy (anonymous):

nope - just didn't get the last stage - guess it's tiredness (0030 here in UK) but I'm not seeing how 3^2y/3^6y+1 = 3^4 ends up being 2y-6y-3=4 and that becomes y=-7/4 .... oh well, thanks for the help folks.

OpenStudy (anonymous):

(9^y)/(27^(2y+1))=81 \[\frac{(3^2)^y}{(3^3)^{(2y+1)}}=3^4\] \[\frac{3^{2y}}{3^{3(2y+1)}}=3^4\] In division (powers of the same base) we subtract the exponents: \[3^{2y-3(2y+1)}=3^4\] Compare the exponents of 3 on both sides of the equation above, and get:: \[2y-3(2y+1)=4\] \[2y-6y-3=4\] \[-4y=7\] \[y=-\frac{7}{4}\]